OpenStudy (anonymous):

What is true about the solutions of a quadratic equation when the radicand of the quadratic formula is zero? Answer No real solutions Two identical rational solutions Two different rational solutions Two irrational solutions

OpenStudy (anonymous):

Two identical rational solutions..

OpenStudy (anonymous):

When D = 0, then two real but equal solutions form..

OpenStudy (klimenkov):

Try to get this using an article http://openstudy.com/study#/updates/5007c20be4b020a2b3bd2e96

OpenStudy (anonymous):

Might not be rational, but will certainly be real. In the case of a quadratic with integer coefficients, it will be rational.

OpenStudy (anonymous):

Example:\[(x-\sqrt2)^2=x^2-2\sqrt2x+2=0 \implies x=\sqrt2\]One (repeated) root, real but not rational.